OpenStudy (anonymous):

A sailboat moves north for a distance of 10.00 km when blown by a wind from the southeast with a force of 5.00 × 10^4 N. How much work was done by the wind? A. 0.00 J B. 3.54 * 10^5 J C. 3.54 * 10^8 J D. 5.00 * 10^8 J

OpenStudy (michele_laino):

here is the situation described by your exercise: 

OpenStudy (anonymous):

yes:)

OpenStudy (michele_laino):

so the requested work is given by the subsequent formula: \[\begin{gathered} W = Fd\cos \left( {45} \right) = 5 \times {10^4} \times {10^3} \times \cos 45 \hfill \\ \hfill \\ \end{gathered} \]